Assuming the market price of coal is p=$300 (inverse supply P=0.2Q), calculate:

A. Quantity supplied

B.MWTA by producers

C.MC of production

D. WTA by producers

E. Total variable costs

F. total revenue

G. producer surplus
Answer #1

(A) Assuming perfect competition, quantity supplied would be 300 = 0.20 or 1500 units.

(B) The supply curve represents the willingness to accept (WTA), and hence, the marginal willingness to accept (MWTA) would be the amount of WTA per unit quantity. The WTA is ap ao or MW.T.A = _(0.2Q ao or MWTA = 0.2 .

(C) The supply curve is determined as P = MC over the average variable cost. In this case, the supply is P = 0.2Q , and hence, the marginal cost will be MC 0.29 = . As the MC is linear, we may say that it is never below AVC for Q>0.

(D) WTA by producers is P = 0.2Q . At current equilibrium quantity of 1500 units, the WTA of producers is the equilibrium price as P 0.26)= 0.2 * 1500 300 .

(E) The marginal cost is MC 0.29 = , and is found by MC--(TCY do , and hence we have TC = int MC mathrm{d}Q or TC = int 0.2Q mathrm{d}Q or TC = 0.2rac{Q^2}{2} + F or TC = 0.1Q^2 + F , where F - the fixed cost, is the integral constant. The total variable cost would be hence TVC = 0.1Q^2 (since TC is TVC+F).

(F) The total revenue would be TR = Q.P or TR = Q(0.2Q) or TR = 0.2Q^2 . In this case, the total revenue would be TR = 0.2(1500)^2 = 450000 dollars.

(G) The cost of producers would be the area below the supply curve as int^{Q^*}_0 P.mathrm{d}Q , where Q* is the equilibrium quantity. Hence, we have int^{Q^*}_0 P.mathrm{d}Q = int^{1500}_0 0.2Q .mathrm{d}Q or int^{Q^*}_0 P.mathrm{d}Q = 0.2 int^{1500}_0 Q .mathrm{d}Q or int^{Q^*}_0 P.mathrm{d}Q = 0.2 [rac{Q^2}{2}]^{1500}_0 or int^{Q^*}_0 P.mathrm{d}Q = 0.2 [rac{1500^2 - 0^2}{2}] or int^{Q^*}_0 P.mathrm{d}Q = 0.1 [1500^2] = 225000 . The producer surplus would be hence total revenue minus this cost, ie PS= 2 500 dollars.